OpenStudy (anonymous):

A particle moving uniformly along the x axis is located at 9.27 m at 1.12 s and at 4.6 m at 3.04 s. Calculate the particle’s average speed during this time interval

OpenStudy (anonymous):

Speed is never negative but velocity can be \[Average speed=\frac{ distance }{ time }\] or \[Average speed=\frac{ \left| final position - initial position \right| }{ \left| time 2 - time 1 \right| }\]

OpenStudy (anonymous):

ohhhhhhh ok. that makes sense. Thank you!

OpenStudy (anonymous):

No problem!